OpenStudy (anonymous):

The Country Porch decided to create their own blend of coffee called Country Blend from a mix of Kona and Columbian beans but they want to minimize their blend cost. The blend will use between 5 pounds and 30 pounds of Kona beans. The blend will use between 7 pounds and 25 pounds of Columbian beans. They will make no more than 40 pounds of Country Blend each week. The Kona beans cost $4.00 per pound. The Columbian beans cost $3.50 per pound. Write the constraints for this scenario.

OpenStudy (anonymous):

5 ≤ x ≤ 30, 7 ≤ y ≤ 25, x+y ≤ 40

OpenStudy (anonymous):

do u need anything else?

OpenStudy (anonymous):

Is the feasible region bounded or unbounded? If unbounded, is there a maximum or minimum? bounded unbounded, minimum unbounded, maximum None of these

OpenStudy (anonymous):

Name the vertices of the feasible region.

OpenStudy (anonymous):

How many pounds of Kona beans should be used in the Country Blend to minimize costs?

OpenStudy (anonymous):

How many pounds of Kona beans should be used in the Country Blend to minimize costs? 5 should be used

OpenStudy (anonymous):

I cant help u with the whole feasible thing

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

do u need help with the whole country porch thing?

OpenStudy (anonymous):

Write the constraints.

OpenStudy (anonymous):

You are about to take a 50-minute test with a mix of computation and word problems. Each computation problem is worth 3 points each. Each word problem is worth 6 points each. You can complete no more than 25 computation problem in 50 minutes. You can complete no more than 10 word problems in 50 minutes. You may answer no more than 15 problems. Let x represent the number of computation problems you answer and y represent the number of word problems you answer. Write the objective function to determine how many of each type of question you need to answer to maximize your score.

OpenStudy (anonymous):

can you please help me find the constraints please

OpenStudy (anonymous):

constraints for what? the country porch question

OpenStudy (anonymous):

You are about to take a 50-minute test with a mix of computation and word problems. Each computation problem is worth 3 points each. Each word problem is worth 6 points each. You can complete no more than 25 computation problem in 50 minutes. You can complete no more than 10 word problems in 50 minutes. You may answer no more than 15 problems. Let x represent the number of computation problems you answer and y represent the number of word problems you answer. Write the objective function to determine how many of each type of question you need to answer to maximize your score. this question

OpenStudy (anonymous):

for the 50min test thing: Assuming your goal is to score maximum points in given 50 minutes of time, the solution is: A: find the function of your equi-scoring curve (say based on time limit) B: find the other limiting function (base on number of problems limit) C: find a point on highest A that also intersects B A: With 50/25 = 2 minutes per comp problem and 50/10 = 5 minutes per word problem you're limited by 2x + 5y <= 50 B: x + y <= 15 You can plot both functions (plot the lines as if <= was = and shade the halfplane under them) for easier interpretation. Assuming x + y = 15 these would meet at 2x + 5(15 - x) = 50 2x + 75 - 5x = 50 -3x = -25 x = 25/3 = 8 and 1/3 which (truncated down) would imply y = 7 Unfortunately that would leave you with 51 minute exam so this is a no-go. You need to reduce either x or y to find the best solution. x = 7 and y = 7 would amount to 7*3 + 7*6 = 63 points and time of 7*2 + 7*5 = 39 minutes which is fine. Reduced y (to 6) would leave you with enough time for 10 comp problems, but unfortunately you can only solve a total of 15 s you have to give up one, so x = 9 and y = 6. In this case you also score 63 points (9*3 + 6*6), but it only takes you 9*2 + 6*5 = 48 minutes to finish the test Reducing y further is pointless as you are limited with 15 problems per exam and you're not using up your quota of 15 as it is (you'd solve another comp problem if it was possible, you have enough time for one more). Reducing x is also pointless as comp problems yield more points per minute of time than word problems so so making time for extra word problems at the cost of comp problems doesn't benefit you at all. Case in point: 8 word problems leave enough time for only 5 comp problems, totaling 5*3 + 8*6 = 63 points, but wasting all 50 minutes of your time (no post-exam talking with that class mate of yours).

OpenStudy (anonymous):

b is the constraits and the rest is the function?

OpenStudy (anonymous):

Im sorry I cant help. This is too confusing! Good luck